### v1

@@ Abstract
We review our current knowledge of thermal and viscous instabilities in accretion discs around compact objects. We begin with classical disc models based on analytic viscosity prescriptions, discussing physical uncertainties and exploring time-dependent solutions of disc evolution. We also review the ionization instability responsible for outbursting dwarf nova and X-ray binary systems, including some detailed comparisons between alpha-based models and the observed characteristics of these systems. We then review modern theoretical work based on ideas around angular momentum transport mediated by magnetic fields, focusing in particular on knowledge gained through local and global computer simulations of MHD processes in discs. We discuss how MHD may alter our understanding of outbursts in white dwarf and X-ray binary systems. Finally, we turn to the putative thermal/viscous instabilities that were predicted to exist in the inner, radiation pressure-dominated regions of black hole and neutron star discs, in apparent contradiction to the observed stability of the high/soft state in black hole X-ray binaries.
@@ 1 Introduction
The seminal 1973 paper by Shakura & Sunyaev [1] laid out a basic model for the radial structure of stationary, geometrically thin accretion discs around black holes. It did so by assuming that the vertically-averaged angular momentum-transporting stress $w$ in the disc was approximately proportional to the vertically-averaged total thermal pressure $P$ in the disc:
@@
One year after the original Shakura & Sunyaev paper, it was pointed by Lightman & Eardley [2] that the radiation pressure and Thomson scattering dominated inner zone was unstable. Assuming vertical hydrostatic and thermal equilibrium, the standard alpha prescription leads to an inverse relationship between vertically integrated stress $W$ and surface mass density $\Sigma$ in the disc, which in turn results in an effectively negative radial mass diffusion coefficient. Fluctuations in surface mass density would then be amplified, likely causing the disc to break up into rings.
@@
In another seminal paper, Shakura & Sunyaev [3] followed this up with an analysis that also accounted for time variations in the thermal balance of the disc, and found that the instability discovered by Lightman & Eardley was part of a more complicated set of instabilities that exist in the inner disc region. At long radial wavelengths $\Lambda$, these instabilities separate into two branches: the anti-diffusion “viscous” instability of Lightman & Eardley, with growth rate $\sim\alpha(H/\Lambda)^{2}\Omega$, and a faster thermal instability with growth rate $\sim\alpha\Omega$. Here $H$ is the local disc vertical thickness, and $\Omega$ is the local orbital angular velocity. The thermal instability arises in the standard alpha prescription because the local heating rate per unit area $Q^{+}\propto\alpha T^{8}/\Sigma$, while the cooling rate $Q^{-}\propto T^{4}/(\kappa\Sigma)$. The temperature dependence of the heating rate at fixed surface mass density is therefore much steeper than that of the cooling rate if Thomson opacity dominates the opacity $\kappa$, so that temperature fluctuations undergo runaway heating or cooling.
@@
Unfortunately, it has never been clear whether these instabilities have a physical reality, or are merely an artifact of the assumptions behind the alpha-prescription. Simple modifications to the alpha prescription can in fact stabilize the radiation pressure dominated inner regions of black hole accretion discs [4]. For example, rapid photon diffusion might decouple turbulent fluctuations from the radiation pressure, so that the stress might scale in proportion to just the pressure in the gas (e.g. [2, 5, 6]). Or perhaps the speeds of turbulent eddies would be limited to the gas sound speed but still be of a size that reflects the radiation pressure scale height, leading to a stress that is proportional to the geometric mean of gas and radiation pressure (e.g. [7]). Both of these prescriptions would stabilize the inner disc. Winds can also help stabilize the disc [8]. Moreover, in the case of discs around supermassive black holes, the opacity might not be dominated by electron scattering, and this could also stabilize the inner regions of the disc [9]. Finally, even if the stress does scale with the total pressure in the disc, that pressure might be dominated by temperature-independent magnetic pressure rather than thermal pressure [10], with dissipation happening above the disc [11], within the disc [12, 13, 14], or a combination of the two [15].
@@
Still another possibility that avoids the thermal and viscous instabilities of the inner regions of black hole accretion discs is that the accretion power that is dissipated at each radius is not radiated away locally, but is instead advected inward. This generates a whole new family of advection-dominated accretion disc solutions based on the Shakura-Sunyaev alpha-prescription that are thermally and viscously stable [16, 17, 18, 19, 20]. The local topology of these solutions and its relation to non-advective models is very nicely summarized in the short paper by Chen et al. [21]. Such models have been applied to quiescent, low/hard, and intermediate states [22, 23] and to near-Eddington accretion states [24] (so-called “slim” accretion discs [16]) of black hole X-ray binaries. However, the high/soft state is still generally modeled as a geometrically thin disc with little advection [22, 24, 25], and the extremely low variability of this state suggests that it is stable [26]. Advective models have also been applied to low luminosity [27] and high luminosity [28] accretion onto supermassive black holes.
@@
We do not discuss models of collimated jets in this review, nor do we discuss so-called MAD (magnetically arrested disc) flows which are an important aspect of the jet problem. These topics are thoroughly discussed in another chapter in this volume. We will discuss magnetocentrifugal winds, however, as they may play an important role in the operation of thermal and viscous instabilities in discs.
@@ 2 Analytic Solutions for Time-Dependent Discs
$\frac{\partial\Sigma}{\partial t}=\frac{1}{r}\,\frac{\partial}{\partial r}\left[\left(\frac{\partial\,\Omega r^{2}}{\partial r}\right)^{-1}\,\frac{\partial}{\partial r}\left(r^{3}\,\frac{\partial\Omega}{\partial r}\,\nu\,\Sigma\right)\right]\,,$ (2)
where $\nu$ is the coefficient of kinematic viscosity [e.g. 2, 29, 3, 30, 31]. The local effective kinematic viscosity in the discs can be associated with various mechanisms that ensure the angular momentum transfer: hydrodynamic [see 32, for a review], magnetorotational turbulence [33], convection-driven turbulence [34] and self-gravity instabilities [35]. In weakly-magnetized accretion discs, where shear and substantial ionization are present, MRI is thought to play a major role, see §3.1.
@@
[…]
@@
In this nonlinear diffusion equation $D$ is the diffusion coefficient:
@@
Eq. (3) follows the view used by Lynden-Bell and Pringle [29] with one notable difference. It assumes a positive value of $F=-g$ for a standard disc, while the torque in its physical definition is negative in the disc. This convention is seen in a number of works and belongs to the scheme when the $\alpha$-ansatz is prescribed in the form (1); it is also consistent with the frequent situation when the index $m$ is non-integer.
To relate $D$ and the kinematic viscosity $\nu$, it is sufficient to express the kinematic viscosity as a product of power law functions
@@
Substitution of (5) into l.h.s. of
@@
Therefore, the disc evolution equation in form (3) has the advantage of allowing one to impose transparent explicit boundary conditions. Also, the flux radiated from one surface of the disc can be written in a unified form applicable to a general case of accretion/decretion discs [38]:
@@
[…]
At large distances around normal or compact stars, where gas pressure dominates and/or opacity is determined by absorption processes in the ionized matter, $D$ is positive and a stable long-term evolution is possible. For $\alpha$-discs, the diffusion coefficient $D$ only weakly depends on the opacity: as a power law with an index of ${1/5}$ or ${1/10}$ for the Thomson or Kramers opacities, respectively. This mitigates the effect of the uncertainty associated with the dependence of the real opacity on the disc parameters. Consequently, $D$ can be regarded as constant in the equation of non-stationary accretion (3).
A method to determine the long-term evolution from the equation of non-stationary accretion (3) depends on the form of the turbulent viscosity coefficient $\nu_{\mathrm{t}}=\nu_{\mathrm{t}}(r,\Sigma)$. If the viscosity $\nu_{\mathrm{t}}$ is a function of the radius only and does not depend on the surface density (i.e. does not depend on time), then $F$ depends linearly on the surface density $\Sigma$. For such cases, when $m=0$ and $a=0$, and Eq. (3) becomes a linear diffusion equation. Note that in this case the characteristic viscous time scale $\tau_{\mathrm{vis}}\sim r^{2}/\nu_{\mathrm{t}}$ is constant in time. In a more general case, $\nu_{\mathrm{t}}$ also depends on the surface density. In particular, if we consider discs with $\alpha$-viscosity, we can represent $\nu_{\mathrm{t}}$ as a power-law function of $\Sigma$ and $r$.
A general solution to the linear diffusion equation can be determined through the expansion in eigenfunctions. The method of superposition enables the formulation of a solution that complies with the specified initial or boundary conditions. In the case of a nonlinear diffusion equation, separation of variables provides solutions only for a certain class of problems. An understanding of the physical properties of problems can, in certain situations, enable the derivation of self-similar solutions. A self-similar solution of the nonlinear diffusion equation describes the evolution when enough time has passed so that the initial state is forgotten.
@@
In the presence of external torques, Eq. (2) is not valid, strictly speaking. If the disc rotates around a binary or in a binary, the tidal interaction exchanges the angular momentum between the disc and the binary orbital motion. In both cases, it is common to assume that the tidal torque is accumulated in the narrow disc ring, the inner or the outer one [39]. Therefore, the influence of a binary on a disc can be represented as an effective boundary condition for Eq. (2).
When accreting onto a black hole, the disc has zero torque at the inner boundary [standard disc by 1]. The other situation — where there is no accretion through the inner boundary, but there is a finite torque at the inner boundary $r_{\mathrm{in}}$ — is thought to occur for rapidly rotating magnetized neutron stars or discs around binary systems. Rafikov [40] showed that there is indeed a whole range of possibilities for the internal boundary condition, with generally non-zero central viscous stress and accretion rates varying from positive to negative values (e.g. if the disc is gaining mass from a central Be-star); see also Nixon and Pringle [41].
The tidal forces of the companion star confine the disc within the Roche lobe of the accretor [42, 43, 44, 45]. Significant perturbations near the outer disc radius lead to the formation and dissipation of shocks and the loss of angular momentum from the disc, which is transferred to the orbital motion. Papaloizou and Pringle [42] showed that the tidal truncation radius is on average $\sim 0.9$ times that of the Roche lobe. This radius is close to that of the last non-intersecting periodic orbit calculated for a three-body problem [43]. Numerical calculations have shown that the tidal stress tensor is only important in a rather narrow ring close to the outer radius. One can therefore choose not to study this region in detail and simply consider it as a $\delta$-type channel, with $r_{\rm out}=\,$const.
@@ 2.1 Linear equation, freely expanding disc
In 1952, Lüst found particular solutions to the linear equation of viscous accretion, proposed by his teacher von Weizsäcker (1948), and described the principles of constructing a general solution to both infinite and finite problems.
Lynden-Bell and Pringle [29] used a method of superposition of particular solutions to the equation of viscous evolution and found general solutions for two types of boundary conditions at the inner boundary: $F=0$ (zero viscous torque at the inner radius, applicable, for example, in the case of accretion onto a black hole) and $\partial F/\partial h=0$ (zero accretion rate). Given the initial distribution $F_{0}(h)$, the solution to the linear differential equation is a linear integral transform:
@@
where $G$ is the Green’s function of Eq. (3) with $m=0$. The problem of solving a linear differential equation with boundary conditions can be viewed as a linear system with an input signal $F_{0}(h_{1})$ and an output signal $F(h,t)$. The Green’s function then acts as the system’s response to a delta impulse input signal $F_{0}(h_{1})=\delta(h_{1}-h_{\rm s})$. The Green’s function itself is obtained as the solution to a boundary value problem with the continuum spectrum of eigenvalues and the Dirac delta function as the initial condition. As a result, $G$ is an integral containing Bessel functions of non-integer order that can be taken analytically by a Hankel transform [see §2.4 in 29].
@@
$\dot{M}(h,t)=\int\limits_{0}^{\infty}G_{\dot{M}}(h,h_{1},t)\,F_{0}(h_{1})\,\mathrm{d}h_{1}~{}/h_{\rm s}\,.$
In the case of the accretion onto a black hole from a viscously spreading ring, the accretion rate at the inner disc boundary has the explicit analytic form [29]:
@@
It is assumed that the ring with mass $M_{\mathrm{disc}}$ was originally located at $r_{s}=h_{s}^{2}/GM$. It is easy to see from (3) and (6) that $D=4\,({l}/{\kappa})^{2}$ for $m=0$.
@@
Pringle [39] found the Green’s functions for an ‘external’ disc with the zero accretion rate at $r_{\mathrm{in}}>0$ and for $\nu\propto r$. In this particular case of $b=1$, the Bessel functions reduce to functions of the form $\sin x/\sqrt{x}$ and $\cos x/\sqrt{x}$. Tanaka [49] found Green’s function solutions for arbitrary $b$ that satisfy either a zero-torque or zero-flux condition when the inner boundary of the disc is located at a finite radius $r_{\mathrm{in}}>0$.
For the zero accretion rate and the finite torque at the inner boundary, $\Sigma$ and $F$ vary with time as $\propto t^{-1+l}$. Keeping in mind that the bolometric flux varies according to (9), we see that the luminosity of such discs decreases slower with time than that of discs with the vanishing inner stress [see also 50, where the numerical solution for a disc around a Kerr black hole and non-vanishing stress is considered in the context of TDE observations]. Near the center, the torque $F$ develops a flat profile, the amplitude of which rises at first and then decreases with time. Rafikov [40] found general late-time asymptotics for non-zero accretion rate and non-zero inner torque (for a general case with $a\neq 0$ too, see §2.3). It was shown that suppression of the central accretion rate always requires stronger central torque. Conversely, the central torque on the disc suppresses mass accretion on the object or even reverses it to an outflow, as in decretion discs around Be-stars.
A general property of a linear differential equation is that the superposition of its solutions yields another valid solution. Using the $\nu\propto r$ case as an example, Nixon and Pringle [41] showed how to obtain the Green’s function in the case of a general finite torque boundary condition at the inner boundary
@@
Balbus [51] has derived a form of the thin disc diffusion equation that is valid for both the Schwarzschild and Kerr geometries. The Green’s function solutions can be calculated in terms of quadratures by using a combination of WKB technique, local analysis and matched asymptotic expansions. Mummery [52] computed analytically the leading order Green’s function solutions (for $F$, $\Sigma$ and $\dot{M}$) [^2] In Mummery [52] $\mu=b-1/2$ and $\alpha=1-b/2$. of the general relativistic thin disc equations for a vanishing stress at the innermost stable circular orbit. It turns out that the peak mass accretion rate across the ISCO is lower for more rapidly rotating Kerr black holes. When the accretion efficiency dependence on the spin is taken into account, the rest-frame luminosity of the more rapidly spinning black holes is the largest.

### v2

@@ Abstract
We review our current knowledge of thermal and viscous instabilities in accretion discs around compact objects. We begin with classical disc models based on analytic viscosity prescriptions, discussing physical uncertainties and exploring time-dependent solutions of disc evolution. We also review the ionization instability responsible for outbursting dwarf nova and X-ray binary systems, including some detailed comparisons between alpha-based models and the observed characteristics of these systems. We then review modern theoretical work based on ideas around angular momentum transport mediated by magnetic fields, focusing in particular on knowledge gained through local and global computer simulations of MHD processes in discs. We discuss how magnetohydrodynamics (MHD) may alter our understanding of outbursts in white dwarf and X-ray binary systems. Finally, we turn to the putative thermal/viscous instabilities that were predicted to exist in the inner, radiation pressure-dominated regions of black hole and neutron star discs, in apparent contradiction to the observed stability of the high/soft state in black hole X-ray binaries.
@@ 1 Introduction
The seminal 1973 paper by sha73 laid out a basic model for the radial structure of stationary, geometrically thin accretion discs around black holes. It did so by assuming that the vertically-averaged angular momentum-transporting stress $w$ in the disc was approximately proportional to the vertically-averaged total thermal pressure $P$ in the disc [^1] The exact definition used by sha73 (equation 1.2 of their paper) actually defined the stress as $\alpha$ times the density times the square of the sound speed, the latter being defined in terms of the total thermal energy density and the density. Apart from factors of order unity, this is equivalent to Eq. (1).:
@@
One year after the original Shakura & Sunyaev paper, it was pointed out by lig74 that the radiation pressure and Thomson scattering dominated inner zone was unstable. Assuming vertical hydrostatic and thermal equilibrium, the standard alpha prescription leads to an inverse relationship between vertically integrated stress $W$ and surface mass density $\Sigma$ in the disc, which in turn results in an effectively negative radial mass diffusion coefficient. Fluctuations in surface mass density would then be amplified, likely causing the disc to break up into rings.
@@
In another seminal paper, sha76 followed this up with an analysis that also accounted for time variations in the thermal balance of the disc, and found that the instability discovered by Lightman & Eardley was part of a more complicated set of instabilities that exist in the inner disc region. At long radial wavelengths $\Lambda$, these instabilities separate into two branches: the anti-diffusion “viscous” instability of Lightman & Eardley, with growth rate $\sim\alpha(H/\Lambda)^{2}\Omega$, and a faster thermal instability with growth rate $\sim\alpha\Omega$. Here $H$ is the local disc vertical thickness, and $\Omega$ is the local orbital angular velocity. The thermal instability arises in the standard alpha prescription because the local heating rate per unit area $Q^{+}\propto\alpha T^{8}/\Sigma$, while the cooling rate $Q^{-}\propto T^{4}/(\kappa\Sigma)$. The temperature dependence of the heating rate at fixed surface mass density is therefore much steeper than that of the cooling rate if Thomson opacity dominates the opacity $\kappa$, causing runaway heating or cooling.
@@
Unfortunately, it has never been clear whether these instabilities have a physical reality, or are merely an artifact of the assumptions behind the alpha-prescription. Simple modifications to the alpha prescription can in fact stabilize the radiation pressure dominated inner regions of black hole accretion discs [pir78]. For example, rapid photon diffusion might decouple turbulent fluctuations from the radiation pressure, so that the stress might scale in proportion to just the pressure in the gas [e.g. lig74, sak81, ste84]. Or perhaps the speeds of turbulent eddies would be limited to the gas sound speed but still be of a size that reflects the radiation pressure scale height, leading to a stress that is proportional to the geometric mean of gas and radiation pressure [e.g. taa84]. Both of these prescriptions would stabilize the inner disc. Winds can also help stabilize the disc [pir77]. Moreover, in the case of discs around supermassive black holes, the opacity might not be dominated by electron scattering, and this could also stabilize the inner regions of the disc [pri76]. Finally, even if the stress does scale with the total pressure in the disc, that pressure might be dominated by temperature-independent magnetic pressure rather than thermal pressure [shi90], with dissipation happening above the disc [fie93], within the disc [par03, beg07, oda09], or a combination of the two [beg15].
@@
Still another possibility that avoids the thermal and viscous instabilities of the inner regions of black hole accretion discs is that the accretion power that is dissipated at each radius is not radiated away locally, but is instead advected inward. This generates a whole new family of advection-dominated accretion disc solutions based on the Shakura-Sunyaev alpha-prescription that are thermally and viscously stable [abr88, nar94, abr95, nar95a, nar95b]. The local topology of these solutions and its relation to non-advective models is very nicely summarized in the short paper by che95. Such models have been applied to quiescent, low/hard, and intermediate states [esi97, esi98] and to near-Eddington accretion states [str11] [so-called “slim” accretion discs, abr88] of black hole X-ray binaries. However, the high/soft state is still generally modeled as a geometrically thin disc with little advection [esi97, str11, dav06], and the extremely low variability of this state suggests that it is stable [gie04]. Advective models have also been applied to low luminosity [nar98] and high luminosity [szu96] accretion onto supermassive black holes.
@@
We do not discuss models of collimated jets in this review, nor do we discuss so-called MAD (magnetically arrested disc) flows which are an important aspect of the jet problem. These topics are thoroughly discussed elsewhere. We will discuss magnetocentrifugal winds, however, as they may play an important role in the operation of thermal and viscous instabilities in discs.
@@ 2 Analytic Solutions for Time-Dependent Discs
In this section we discuss a variety of analytic solutions for the time-dependent behavior of accretion discs. Construction of such solutions requires various simplifying assumptions, particularly in the stress prescription (or equivalently, an effective viscosity). Although real accretion disks may not fully satisfy these assumptions, such solutions are nevertheless valuable in understanding the behavior of more complex simulations with more realistic physics, particularly MHD. Moreover, the present-day limitations of computational power mean that analytic solutions are still the only way to model the evolution of discs on viscous time scales. Eventually, one hopes that simulation data can be used to improve upon the assumptions that are used in constructing these solutions.
@@
$\frac{\partial\Sigma}{\partial t}=\frac{1}{r}\,\frac{\partial}{\partial r}\left[\left(\frac{\partial\,\Omega r^{2}}{\partial r}\right)^{-1}\,\frac{\partial}{\partial r}\left(-r^{3}\,\frac{\partial\Omega}{\partial r}\,\nu\,\Sigma\right)\right]\,$ (2)
[e.g. lig74, lyn-pri1974, sha76, Pringle1981, lin-pringle1987, shakura_etal2018], where $\nu$ is the coefficient of kinematic viscosity. The local effective kinematic viscosity in the discs can be associated with various mechanisms that ensure the angular momentum transfer: hydrodynamic [see Lesur+2023, for a review], magnetorotational turbulence [bal91], convection-driven turbulence [Lin-Papaloizou1980] and self-gravity instabilities [Paczynski1978]. In weakly-magnetized accretion discs, where shear and substantial ionization are present, magnetorotational (MRI) turbulence is thought to play a major role, see §3.1.
@@
Mathematically, Eq. (2) is a parabolic equation of the second order in partial derivatives, that is, a diffusion equation for $\Sigma$, which generally is nonlinear because the viscosity can depend on the surface density [lig74]. In the case of Keplerian orbits, the equation can be written much more simply. In particular, if it is written in terms of the $z$-integrated viscous torque $F=-r\,2\pi r\Sigma\nu r\,{\partial\Omega}/{\partial r}$, which is a couple exerted on a ring of material by the disc interior to $r$ and becomes $F=3\,\pi\,h\,\nu\,\Sigma$, and the specific Keplerian angular momentum $h=\Omega_{\scriptscriptstyle{\mathrm{K}}}\,r^{2}=\sqrt{GMr}$. The disc-evolution equation can then be written as [filipov1984, lyub-shak1987]:
@@
In this nonlinear diffusion equation $m$ and $n$ are the scaling exponents and $D$ is the diffusion coefficient, which follows from imposing local thermal equilibrium:
@@
Eq. (3) follows the view used by lyn-pri1974 with their $g(h,t)=F(h,t)$. It assumes a positive value of the viscous torque $F$ for a standard disc, while the net viscous torque (that is, the difference of the outer and inner torques, applied to a ring) is negative in the disc. This convention is seen in a number of works and is consistent with the frequent situation when the parameter $m$ is non-integer.
To relate $D$ to the kinematic viscosity $\nu$, it is sufficient to express the latter as a product of power law functions
@@
Here $\nu_{0}$ is a dimensional normalization factor. Substitution of (5) into the l.h.s. of
@@
The normalization $\nu_{0}$ can be predicted using (6) if the vertical structure of the disc is calculated, so that $D$ is found via Eq. (4).
@@
Therefore, the disc evolution equation in form (3) has the advantage of allowing one to impose transparent explicit boundary conditions. Also, the flux radiated from one surface of the disc can be written in a unified form applicable to a general case of accretion/decretion discs [rafikov2013]:
@@
The characteristics of viscosity and the vertical structure of the disc determine the parameters $a$ and $b$, as well as $D$, $m$ and $n$; see Table 2 below [see, e.g., shakura_etal2018]. Usually the viscosity depends on the disc physical conditions and thus on the accretion rate in the disc. For example, the anti-diffusion viscous instability found by lig74 in the inner radiation-pressure dominated and Thomson scattering zone corresponds to the negative effective diffusion coefficient $D$ with a=2a=-2 in Eq. (5) [^2] The solution to the angular momentum balance equation (8) is $\dot{M}\,h\,f(r)=F$, where $f(r)$ is some function. In the $\alpha$-disc, the viscous torque $F=2\pi r^{2}\,2z\,\alpha\,P$. Substituting the pressure from the hydrostatic balance $P\sim\Sigma\,z\Omega_{\scriptscriptstyle{\mathrm{K}}}^{2}$ in the latter, one obtains $\dot{M}\,f(r)\sim 4\pi\,\alpha z^{2}\,\Sigma\Omega_{\scriptscriptstyle{\mathrm{K}}}$. In the radiation pressure dominated zone with Thomson opacity, the half-thickness of the disc $z\sim 3\dot{M}\sigma_{\rm T}/(8\pi c)$, which yields $\Sigma\propto\dot{M}^{-1}$. Equating the viscous torque to the expression in (5), we find the dependence for the kinematic turbulent viscosity: $\nu\propto\Sigma^{-2}$..
At large distances around normal or compact stars, where gas pressure dominates and/or opacity is determined by absorption processes in the ionized matter, $D$ is positive and a stable long-term evolution is possible. For $\alpha$-discs, the diffusion coefficient $D$ depends only weakly on the opacity: as a power law with an index of 1/5{1/5} or 1/10{1/10} for Thomson or Kramers opacities, respectively. This mitigates the effect of the uncertainty associated with the dependence of the real opacity on the disc parameters. Consequently, $D$ can be regarded as constant in the equation of non-stationary accretion (3).
A method to determine the long-term evolution from the equation of non-stationary accretion (3) depends on the form of the turbulent viscosity coefficient $\nu=\nu(r,\Sigma)$. If the viscosity $\nu$ is a function of the radius only and does not depend on the surface density (i.e. does not depend on time), then $F$ depends linearly on the surface density $\Sigma$. For such cases, when $m=0$ and $a=0$, Eq. (3) becomes a linear diffusion equation. Note that in this case the characteristic viscous time scale $\tau_{\mathrm{vis}}\sim r^{2}/\nu$ is constant in time. In a more general case, $\nu$ also depends on the surface density. In particular, if we consider discs with $\alpha$-viscosity, we can represent $\nu$ as a power-law function of $\Sigma$ and $r$.
A general solution to the linear diffusion equation can be determined through an eigenfunction expansion. The method of superposition enables the formulation of a solution that complies with the specified initial or boundary conditions. In the case of a nonlinear diffusion equation, separation of variables provides solutions only for a certain class of problems. An understanding of the physical properties of problems can, in certain situations, enable the derivation of self-similar solutions. A self-similar solution of the nonlinear diffusion equation describes the evolution when enough time has passed so that the initial state is forgotten.
@@
In the presence of external torques, Eq. (2) is not valid, strictly speaking. If the disc rotates around a binary or in a binary, the tidal interaction exchanges the angular momentum between the disc and the binary orbital motion. In both cases, it is common to assume that the tidal torque is accumulated in a narrow disc ring, the inner or the outer one [pringle1991]. Therefore, the influence of a binary on a disc can be represented as an effective boundary condition for Eq. (2).
When accreting onto a black hole, the disc is often assumed to have zero torque at the inner boundary [standard disc by sha73]. The other situation — where there is no accretion through the inner boundary, but there is a finite torque at the inner boundary $r_{\mathrm{in}}$ — is thought to occur for rapidly rotating magnetized neutron stars or discs around binary systems. rafikov2016 showed that there is indeed a whole range of possibilities for the internal boundary condition, with generally non-zero central viscous stress and accretion rates varying from positive to negative values (e.g. if the disc is gaining mass from a central Be-star); see also Nixon-Pringle2021.
The tidal forces of a companion star confine the disc within the Roche lobe of the accretor [pap-pri1977, paczynski1977, ich-osa1994, hameury-lasota2005]. Significant perturbations near the outer disc radius lead to the formation and dissipation of shocks and the loss of angular momentum from the disc, which is transferred to the orbital motion. pap-pri1977 showed that the tidal truncation radius is on average $\sim 0.9$ times that of the Roche lobe. This radius is close to that of the last non-intersecting periodic orbit calculated for a three-body problem [paczynski1977]. Numerical calculations have shown that the tidal stress tensor is only important in a rather narrow ring close to the outer radius. Therefore, one can choose to not study this region in detail by considering tidal interactions to occur in a zero-width region and assuming that $r_{\rm out}$ is constant.
@@ 2.1 Linear equation, freely expanding disc
In 1952, lust1952 found particular solutions to the linear equation of viscous accretion, proposed by his teacher von Weizsäcker (weizsaecker1948), and described the principles of constructing a general solution to the evolution of discs with either infinite or finite geometrical extent.
lyn-pri1974 used a method of superposition of particular solutions to the equation of viscous evolution and found general solutions for two types of boundary conditions at the inner boundary: $F=0$ (zero viscous torque at the inner radius, applicable, perhaps, to the case of accretion onto a black hole) and $\partial F/\partial h=0$ (zero accretion rate). Given the initial distribution $F_{0}(h)$, the solution to the linear differential equation is a linear integral transform:
@@
where $G$ is the Green’s function of Eq. (3) with $m=0$ ($a=0$). The problem of solving a linear differential equation with boundary conditions can be viewed as a linear system with an input signal $F_{0}(h_{1})$ and an output signal $F(h,t)$. The Green’s function then acts as the system’s response to a delta impulse input signal $F_{0}(h_{1})=\delta(h_{1}-h_{\rm s})$. The Green’s function itself is obtained as the solution to a boundary value problem with the continuum spectrum of eigenvalues and the Dirac delta function as the initial condition. As a result, $G$ is an integral containing Bessel functions of non-integer order that can be taken analytically by a Hankel transform [see §2.4 in lyn-pri1974]. An example of the Green’s function $G(h,1,t)$ is shown in Fig. 1.
@@
$\dot{M}(h,t)=\int\limits_{0}^{\infty}G_{\dot{M}}(h,h_{1},t)\,F_{0}(h_{1})\,\mathrm{d}h_{1}/h_{\rm s}\,.$
In the case of the accretion onto a black hole from a viscously spreading ring, the accretion rate at the inner disc boundary has the explicit analytic form [lyn-pri1974]:
@@
It is assumed that the ring with mass $M_{\mathrm{disc}}$ was originally located at $r_{s}=h_{s}^{2}/GM$. It is easy to see from (3) and (6) that $D=4\,({l}/{\kappa})^{2}$ for $m=0$. Here, $l>0$ and $b<2$.
@@
Here $\nu_{\mathrm{s}}=\nu_{0}\,r_{\mathrm{s}}^{b}$. Assuming the values of $l$, original radius $r_{s}$, and time $t_{\mathrm{max}}$ are known, the normalization factor $\nu_{0}$ can be obtained.
@@
pringle1991 found the Green’s functions for an ‘external’ disc with zero accretion rate at $r_{\mathrm{in}}>0$ and for $\nu\propto r$. In this particular case of $b=1$, the Bessel functions reduce to functions of the form $\sin x/\sqrt{x}$ and $\cos x/\sqrt{x}$. tanaka2011 found Green’s function solutions for arbitrary $b$ that satisfy either a zero-torque or zero-flux condition when the inner boundary of the disc is located at a finite radius $r_{\mathrm{in}}>0$.
For zero accretion rate and a finite torque at the inner boundary, $\Sigma$ and $F$ vary with time as $\propto t^{-1+l}$. Keeping in mind that the bolometric flux varies according to (9), we see that the luminosity of such discs decreases more slowly with time than that of discs with vanishing inner stress [see also Balbus-Mummery2018, where the numerical solution for a disc around a Kerr black hole and non-vanishing stress is considered in the context of TDE observations]. Near the center, the torque $F$ develops a flat profile, the amplitude of which rises at first and then decreases with time. rafikov2016 found general late-time asymptotics for non-zero accretion rate and non-zero inner torque (for a general case with $a\neq 0$ too, see §2.3). It was shown that suppression of the central accretion rate always requires stronger central torque. Conversely, the central torque on the disc suppresses mass accretion on the object or even reverses it to an outflow, as in decretion discs around Be-stars.
A general property of a linear differential equation is that the superposition of its solutions yields another valid solution. Using the $\nu\propto r$ case as an example, Nixon-Pringle2021 showed how to obtain the Green’s function in the case of a general finite torque boundary condition at the inner boundary
@@
Figure 1: Green function $G(h,1,t)$ versus specific angular momentum $h$ for freely expanding disc (left) and for disc with the outer radius fixed so that $h_{\rm out}=2$. Thicker curves correspond to later times. For the bounded disc, the condition of zero accretion rate is applied at the outer radius.
Balbus2017 has derived a form of the thin disc diffusion equation that is valid for both the Schwarzschild and Kerr geometries. The Green’s function solutions can be calculated in terms of quadratures by using a combination of WKB technique, local analysis and matched asymptotic expansions. Mummery2023 computed analytically the leading order Green’s function solutions (for $F$, $\Sigma$ and $\dot{M}$) [^3] In Mummery2023, $\mu=b-1/2$ and $\alpha=1-b/2$. of the general relativistic thin disc equations for a vanishing stress at the innermost stable circular orbit. It turns out that the peak mass accretion rate across the innermost stable circular orbit (ISCO) is lower for more rapidly rotating Kerr black holes. When the accretion efficiency dependence on the spin is taken into account, the rest-frame luminosity of the more rapidly spinning black holes is the largest.